Solve 7x+9y=3


5x+7y=1 using matrix method

1
Expert's answer
2021-12-08T11:06:26-0500

7x+9y=3

5x+7y=1

write the equation in matrix notation

"\begin{bmatrix}
 7 & 9 \\
 5 & 7
\end{bmatrix}" * "\begin{bmatrix}
 x \\
 y
\end{bmatrix}" ="\begin{bmatrix}
 3 \\
 1
\end{bmatrix}"

"X=A^{-1}B"

where "X=\begin{bmatrix}
 x \\
 y
\end{bmatrix}"

"A^{-1}=\frac{1}{det}\begin{bmatrix}
 7 & -9 \\
 -5 & 7
\end{bmatrix}"

"B=\begin{bmatrix}
 3 \\
 1
\end{bmatrix}"

det=(7*7)-(5*9)

=4

"\begin{bmatrix}
 x\\
 y
\end{bmatrix}=\frac 1 4*\begin{bmatrix}
 7& -9\\
 -5 & 7
\end{bmatrix}* \begin{bmatrix}
 3 \\
 1
\end{bmatrix}"


"\begin{bmatrix}
 x\\
 y
\end{bmatrix}=\frac 1 4*\begin{bmatrix}
 12\\
 -8
\end{bmatrix}"

"\begin{bmatrix}
 x\\
 y
\end{bmatrix}=\begin{bmatrix}
 3\\
 -2
\end{bmatrix}"

x=3 and y=-2